OpenStudy (anonymous):

Please Help A child has nine cards numbered 1 through 9. The child places three cards in a row to form a three-digit number. Find the probability that the number is larger than 800.

OpenStudy (anonymous):

to be greater than 800, the first digit must be either 8 or 9. Because cause there is no 0, the second digit can be any of the 8 numbers remaining, and the third can be any of the 7 numbers remaining. So, P = (2 * 8 * 7) / (9 * 8 * 7) = 1/4